OpenStudy (anonymous):

Johnny read an entire series of books with a total of 4,416 pages. In the equation below, d is the number of days it took Johnny to read the book series. 4,416 = 24d What is the unit rate in the equation?

OpenStudy (anonymous):

The unit rate is \(d\): \[4416=24d~~\Rightarrow~~\frac{4416}{24}=\frac{24d}{24}~~\Rightarrow~~d=\frac{4416}{24}=184\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw